bralyn:

You and 2 friends have a job cleaning houses. You split the total money you make so that you each get the same amount. On the first day, you earn $93. The second day, you earn $75. The third day, you earn $108 . How much money do you each get for 3 days of work? CLEAR CHECK $92 per person $204 per person $214 per person $276 per person

bralyn:

@Joe348

bralyn:

@MidNiGhtSkY

bralyn:

@snowflake0531

AestheticValFrand:

why are we putting @ imma go sorry

snowflake0531:

First, add 93+75+108

QuestionCoveBot:

Step 1: Add all the numbers up. 75+93+108= What does that equal then?

AestheticValFrand:

nvm imma help

bralyn:

im bck

bralyn:

276

QuestionCoveBot:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bralyn 276 \(\color{#0cbb34}{\text{End of Quote}}\) Good. Now divide by 2. 276/2

AestheticValFrand:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @QuestionCoveBot Step 1: Add all the numbers up. 75+93+108= What does that equal then? \(\color{#0cbb34}{\text{End of Quote}}\) well 75+ 93+ 108 = 276

bralyn:

somebody else answer that

QuestionCoveBot:

Also, you divide by 2 because there is two people. Forgot that part. .-.

AestheticValFrand:

oh sorry

bralyn:

so whats the anser

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @QuestionCoveBot Also, you divide by 2 because there is two people. Forgot that part. .-. \(\color{#0cbb34}{\text{End of Quote}}\) It's divide by 3 you and 2 other friends, 3 ppl

AestheticValFrand:

it equals 276/2

bralyn:

somebody else anser it i ansered the other one

QuestionCoveBot:

Oop- divide by 3, sorry about that. ._. It would look like 276/3

bralyn:

and im not good a dividing

AestheticValFrand:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bralyn and im not good a dividing \(\color{#0cbb34}{\text{End of Quote}}\) well it is 276/3 !

bralyn:

92 dng

QuestionCoveBot:

Good job! :)

AestheticValFrand:

Good Job!!!